# SkillRT: Compiling Skills for Efficient Execution Everywhere

### 03 Apr 2026 v1

@@ 6.6. Evaluating Environment Binding
The preceding two experiments conducted testing in fully provisioned environments. However, in real-world skill scenarios, missing environment dependencies during agent execution are commonplace. Resolving such environmental issues presents significant challenges for weaker LLMs, frequently resulting in skill execution failures. SkillRT addresses this through an environment binding mechanism in its AOT compilation phase. This mechanism shifts environment dependency installation from the LLM execution phase to pre-execution preprocessing, allowing the LLM to focus solely on skill logic.
@@ 6.7. Evaluating Concurrency Extraction
SkillRT extracts more parallelism opportunities within skills and improves efficiency. We evaluate three parallelism types: data-level parallelism (DLP), instruction-level parallelism (ILP), and thread-level parallelism (TLP). DLP applies when instructions within a single skill batch-process large amounts of independent data. ILP applies when multiple independent instructions or code segments within a skill can execute concurrently. TLP applies when multiple independent sub-agents run within a skill, each processing self-contained subtasks without cross-dependencies. Figure 16 demonstrates SkillRT’s performance improvements across these three parallelism types.
@@ 6.8. Evaluating Code Solidification
If solidified code execution causes task failure, SkillRT’s safety mechanism triggers fallback, re-enabling LLM code generation to ensure correctness.
@@ 7. Discussion
Non-determinism in skill compilation. Unlike traditional program compilation, skill compilation takes natural language as input, which inherently introduces some non-determinism into the compilation process (wang2023selfconsistency, ). However, the LLM executing skills differs fundamentally from traditional hardware such as CPUs (hennessy2019architecture, ), and possesses an inherent tolerance for input variability that enables skills to execute correctly. Moreover, SkillRT’s rigorous compilation optimization passes and rollback mechanisms ensure that compiled skills consistently deliver stable performance improvements across most downstream tasks.
@@ 8. Conclusion
Skills have emerged as a new code form in the agent era. However, after analyzing over 100,000 skills, we discover a significant mismatch between skills and underlying LLMs. To address this, we propose SkillRT, a compilation and runtime system specifically tailored for skills. SkillRT leverages compilation techniques such as AOT and JIT to optimize skill structure, extract environment dependencies, and identify parallelism opportunities. As models and agent harnesses continue to diversify, skill compilation provides a systematic pathway to transform skills into portable components, rather than fragile prompts.

### 06 Apr 2026 v2

@@ 6.6. Evaluating Environment Binding
The preceding two experiments conducted testing in fully provisioned environments. However, in real-world skill scenarios, missing environment dependencies during agent execution are commonplace. Resolving such environmental issues presents significant challenges for weaker LLMs, frequently resulting in skill execution failures. SkVM addresses this through an environment binding mechanism in its AOT compilation phase. This mechanism shifts environment dependency installation from the LLM execution phase to pre-execution preprocessing, allowing the LLM to focus solely on skill logic.
@@ 6.7. Evaluating Concurrency Extraction
SkVM extracts more parallelism opportunities within skills and improves efficiency. We evaluate three parallelism types: data-level parallelism (DLP), instruction-level parallelism (ILP), and thread-level parallelism (TLP). DLP applies when instructions within a single skill batch-process large amounts of independent data. ILP applies when multiple independent instructions or code segments within a skill can execute concurrently. TLP applies when multiple independent sub-agents run within a skill, each processing self-contained subtasks without cross-dependencies. Figure 16 demonstrates SkVM’s performance improvements across these three parallelism types.
@@ 6.8. Evaluating Code Solidification
If solidified code execution causes task failure, SkVM’s safety mechanism triggers fallback, re-enabling LLM code generation to ensure correctness.
@@ 7. Discussion
Non-determinism in skill compilation. Unlike traditional program compilation, skill compilation takes natural language as input, which inherently introduces some non-determinism into the compilation process (wang2023selfconsistency, ). However, the LLM executing skills differs fundamentally from traditional hardware such as CPUs (hennessy2019architecture, ), and possesses an inherent tolerance for input variability that enables skills to execute correctly. Moreover, SkVM’s rigorous compilation optimization passes and rollback mechanisms ensure that compiled skills consistently deliver stable performance improvements across most downstream tasks.
@@ 8. Conclusion
Skills have emerged as a new code form in the agent era. However, after analyzing over 100,000 skills, we discover a significant mismatch between skills and underlying LLMs. To address this, we propose SkVM, a compilation and runtime system specifically tailored for skills. SkVM leverages compilation techniques such as AOT and JIT to optimize skill structure, extract environment dependencies, and identify parallelism opportunities. As models and agent harnesses continue to diversify, skill compilation provides a systematic pathway to transform skills into portable components, rather than fragile prompts.